OpenStudy (anonymous):

The area of a parallelogram is 112.34 square inches and the base is 13.7 inches. What is the height of this parallelogram?

OpenStudy (misssunshinexxoxo):

Do you know formula for a parallelogram?

OpenStudy (anonymous):

no

OpenStudy (misssunshinexxoxo):

What is 112.34/13.7?

OpenStudy (anonymous):

8.2

OpenStudy (anonymous):

11.9

OpenStudy (misssunshinexxoxo):

8.2 is your answer

OpenStudy (anonymous):

how

OpenStudy (anonymous):

HOWWWWWW

OpenStudy (misssunshinexxoxo):

What's 13.7 x 8.2?

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

\[\large Area~=~ Base \times Height\\\large 112.34~=~ 13.7 \times Height\\\large \frac{112.34}{13.7}~=~ Height\]